# Fractional binomial distributions induced by the generalized binomial theorem and their applications

### v1

@@ 1.4. Notation
  • We set =$\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
@@
  • For $x\in\mathbb{R}$, we denote by $\lfloor x\rfloor$ and $\lceil x\rceil$ the greatest integer less than or equal to $x$, and the smallest integer greater than or equal to $x$, respectively.
  • For $a\in\mathbb{R}$, we denote by $\mathrm{sgn}\,a$ the sign of $a$, that is,
@@
  • A positive constant whose concrete expression is not important is denoted by $C_{i,j}$ when it is the $j$-th positive constant in Section $i$
2. The fractional binomial distribution and its properties
2.1. A fractional-order Taylor series
@@
If $\alpha/2\in\mathbb{N}$, (2.1) and (2.2) still hold by regarding the second term of the right-hand side of (2.1) and the right-hand side of (2.2) as 00.
@@
2.2. Moments of the fractional binomial distribution
Suppose that $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. We aim to find concrete expressions of moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ by applying the fractional Taylor series (2.1) and (2.2) to a certain holomorphic function. Let $c\in\mathbb{R}$ and define a differential operator $\mathscr{D}_{c}$ acting on holomorphic functions $f$ on a domain of $\mathbb{C}$ by
@@ 2.2. Moments of the fractional binomial distribution
for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq-1\}$ if $\alpha n\notin\mathbb{N}$ and $z\in\mathbb{C}$ if $\alpha n\in\mathbb{N}$. Moreover, each $W_{c}(m,k)$ is a polynomial in $c$ of degree $m-k$ with integer coefficients (see also Table 1).
@@ 4 $c^{4}$ $14c^{3}+6c^{2}+4c+$ 1 $76c^{2}+12c+$ 7 $64c+$ 6 1
$\hskip 28.45274pt+c\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
@@
$\hskip 28.45274pt+W_{c}(m,m)(\alpha n)_{m+1}z^{m+1}(1+z)^{\alpha n-m-1}+cW_{c}(m,0)(1+z)^{\alpha n}.$
@@
Let $c=0$. Then, the recurrence formula (2.4) gives the Stirling number of the second kind $W_{0}(m,k)=S(m,k)$, which is the number of ways to partition a set of $m$ objects into $k$ non-empty subsets. Therefore, the sequence $\{W_{c}(m,k)\}$ is regarded as an extension of the Stirling numbers of the second kind, which in combinatorics are referred to as the ($c$-)Whitney numbers of the second kind. To the best of our knowledge, this notion was first introduced in [Mez10], and indeed the theorem below tells us that $W_{c}(m,k)$ appears when we express the moments of the fractional binomial distribution.
@@
$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.6)
Here, if $\alpha\in\mathbb{N}$, then the third and fourth terms of the right-hand side are regarded as 00.
@@
Step 1. Let $f$ be a continuous function on $\overline{D}$ such that $f$ is holomorphic on $D$ and $f(-1)=0$. Suppose that a function $g$ defined by
@@
is extended as a continuous function on $\overline{D}$ and the function $(-1/2,1/2)\ni x\longmapsto g(e^{2\pi ix})$ is of class $C^{1}$. Then, we have
@@
In the last equality, we used $f(-1)=0$. Because $m\leq\alpha n$ or $\alpha n\in\mathbb{N}$ holds, we can take the function $h_{c}^{(k)}$, $k=0,1,\dots,m-1$ defined by (2.3) as $f$. Therefore, we obtain
@@
for $\xi\in\mathbb{R}$, from (2.7) and Proposition 2.2.
@@
by using (2.10) with $c$ replaced by $-(c+\alpha n)$, we apply it to (2.9) to obtain
@@
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha\lambda^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.11)
We put $\lambda=x/(1-x)$, $20<x<1/$ 2 and multiply both sides in (2.11) by $(1-x)^{\alpha n}$. Then, we obtain
@@
$\hskip 28.45274pt-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.12)
@@
are extended to holomorphic functions in $x$ on a complex domain containing the set $\{x\in\mathbb{R}\mid 0<x<1\}$ and do not vanish if $\alpha\notin\mathbb{N}$, (2.12) also holds for $10<x<$ 1 by analytic continuation. Moreover, by applying (2.5) with $z=x\omega/(1-x)$ and $\omega\in K_{\alpha}$ to the first term of the right-hand side of (2.12), we have
@@
As seen from (1.7), the normalized constant $Z_{\alpha,x}^{(n)}$ satisfies
$Z_{\alpha,x}^{(n)}=1+O(e^{-\delta n}),\qquad\frac{1}{Z_{\alpha,x}^{(n)}}=1+O(e^{-\delta n})$ (2.13)
as $n\to\infty$ for some $\delta>0$, where $\delta$ can be chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\alpha\in(0,\infty)$. In what follows, we use the notation $O(e^{-\delta n})$ in this sense. Then, we also have the following.
@@
For each $m\in\mathbb{N}_{0}$, there exists some $\delta>0$ such that
@@
Because the left-hand sides of (2.15) and (2.16) are regarded as 00 if $\alpha\in\mathbb{N}$, we may suppose that $\alpha$ is not an integer. It is sufficient to show (2.15) because (2.16) is immediately obtained by replacing $x$ in (2.15) with $1-x$. We have
@@
By the change of variable $s=((1-x)t/x)^{\alpha}$ and [GR15, Equation 12, page 329], the right-hand side of (2.17) becomes
@@
as $n\to\infty$ for some $\delta>0$. Furthermore, Lemma 2.3 implies that
@@
as $n\to\infty$ for some $\delta>0$, by using (2.19) and (2.20). Then, (2.21), (2.22) and Lemma 2.7 imply that the sum of the third and fourth terms on the right-hand side of (2.6) is dominated by
@@
which is $(|c|^{m}+1)O(e^{-\delta n})$ as well. Note that these items are regarded as 0 if $\alpha\in\mathbb{N}$. ∎
Note that Proposition 2.6 with $c=-\alpha nx$ is often used in the rest of this paper. When $c=-\alpha nx$, the second term in the right-hand side of (2.14) is $O(e^{-\delta n})$ with possibly different $\delta>0$. A direct calculation also gives
@@
Let $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is given by $\mu_{\alpha,x}^{(n)}$. Then, for some $\delta>0$, we have
@@
Applying Proposition 2.6 with $c=0$ and $m=1$, we have
@@
as $n\to\infty$, which implies (2.24). We also apply (2.13), (2.23), and Proposition 2.6 with $c=-\alpha nx$ and $m=2$ to obtain
@@
2.3. Uniform bound of $Z_{\alpha,x}^{(n)}$
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. We will give a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to $x$ and $n$, and again we note that $Z_{1,x}^{(n)}=1$.

### v3

@@ 1.5 Notation
  • We set $\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
@@
  • For $\in\mathbb{R}, x\lfloor x\rfloor$ and $\lceil x\rceil$ denote the greatest integer less than or equal to $x$, and the smallest integer greater than or equal to $x$, respectively.
  • For $a\in\mathbb{R}$, $\mathrm{sgn}\,a$ denotes the sign of $a$, that is,
@@
  • A positive constant whose concrete expression is not important is denoted by $C_{i,j}$ when it is the $j$-th positive constant in Section $i$ .
2 The fractional binomial distribution and its properties
2.1 A fractional-order Taylor series
@@
If $\alpha/2\in\mathbb{N}$, (2.1) and (2.2) still hold by regarding the second term of the right-hand side of (2.1) and the right-hand side of (2.2) as 0.
@@
2.2 Moments of the fractional binomial distribution
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. We now derive the explicit expressions for the moments of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ by applying Proposition 2.1 to a certain holomorphic function. For $c\in\mathbb{R}$, we define a differential operator $\mathscr{D}_{c}$ acting on holomorphic functions $f$ on a domain of $\mathbb{C}$ by
@@ 2.2 Moments of the fractional binomial distribution
for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq-1\}$ if $\alpha n\notin\mathbb{N}$ and $z\in\mathbb{C}$ if $\alpha n\in\mathbb{N}$. Moreover, each $W_{c}(m,k)$ is a polynomial in $c$ of degree mkm-k with integer coefficients.
See Table 1 for concrete expressions of some of $W_{c}(m,k)$.
Table 1: Expressions of $W_{c}(m,k)$ for $m=0,1,2,3,4$.
@@ 4 $c^{4}$ $4c^{3}+6c^{2}+4c+1$ $6c^{2}+12c+7$ $4c+6$ 1
$\quad+c\sum_{k=0}^{m}W_{c}(m,k)(\alpha n)_{k}z^{k}(1+z)^{\alpha n-k}$
@@
$\quad+W_{c}(m,m)(\alpha n)_{m+1}z^{m+1}(1+z)^{\alpha n-m-1}+cW_{c}(m,0)(1+z)^{\alpha n}.$
@@
Let $c=0$. Then, the recurrence formula (2.4) gives the Stirling number of the second kind $W_{0}(m,k)=S(m,k)$, which is the number of ways to partition a set of $m$ objects into $k$ non-empty subsets. Therefore, the sequence $\{W_{c}(m,k)\}$ is regarded as an extension of the Stirling numbers of the second kind, which in combinatorics are referred to as the ($c$-)Whitney numbers of the second kind. To the best of our knowledge, this notion was first introduced in [25], and indeed the theorem below tells us that $W_{c}(m,k)$ appears when we express the moments of the fractional binomial distribution.
@@
$\quad-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\quad-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.6)
Here, if $\alpha\in\mathbb{N}$, then the third and fourth terms of the right-hand side are regarded as 0.
@@
Step 1. Let $f$ be a continuous function on $\overline{D}$ such that $f$ is holomorphic on $D$ and f(1)=0f(-1)=0. Suppose that a function $g$ defined by
@@
is extended as a continuous function on $\overline{D}$ and the function $(-1/2,1/2)\ni x\longmapsto f(e^{2\pi ix})$ is of class $C^{1}$. Then, we have
@@
In the last equality, we used f(1)=0f(-1)=0. Because $m\leq\alpha n$ or $\alpha n\in\mathbb{N}$ holds, we can take the function $h_{c}^{(k)}$, $k=0,1,\dots,m-1$ defined by (2.3) as $f$. Therefore, we obtain
@@
for $\xi\in\mathbb{R}$, from (2.7) and Proposition $2.$
@@
by using (2.10) with $c$ replaced by $(c+\alpha n)$, we apply it to (2.9) to obtain
@@
$\quad-(-1)^{m}\frac{\alpha\lambda^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.11)
We put $\lambda=x/(1-x)$, $0<x<1/2$ and multiply both sides in (2.11) by $(1-x)^{\alpha n}$. Then, we obtain
@@
$\quad-\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{c}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\quad-(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-(c+\alpha n)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t.$ (2.12)
@@
are extended to holomorphic functions in $x$ on a complex domain containing the set $\{x\in\mathbb{R}\mid 0<x<1\}$ and do not vanish if $\alpha\notin\mathbb{N}$, (2.12) also holds for $0<x<1$ by the identity theorem. Moreover, by applying (2.5) with $z=x\omega/(1-x)$ and $\omega\in K_{\alpha}$ to the first term of the right-hand side of (2.12), we have
@@
  1. (1)
    As $n\to\infty$,
    $Z_{\alpha,x}^{(n)}=1+O(e^{-\delta n}),\qquad\frac{1}{Z_{\alpha,x}^{(n)}}=1+O(e^{-\delta n}).$ (2.13)
  2. (2)
    For each $m\in\mathbb{N}_{0}$,
@@
Here, the symbol $O(e^{-\delta n})$ means that there exist positive constants $C$ and $\delta$ depending only on $x$, $\alpha$ (and $m$) that are chosen uniformly on any compact subset of $x\in(0,1)$ and any compact subset of $\alpha\in(0,\infty)$ such that the modulus of the corresponding term is dominated by $Ce^{-\delta n}$.
In what follows, we use the notation $O(e^{-\delta n})$ in this sense; $\delta$ may change from line to line.
@@
Because the left-hand sides of (2.15) and (2.16) are regarded as 0 if $\alpha\in\mathbb{N}$, we may suppose that $\alpha$ is not an integer. It is sufficient to show (2.15) because (2.16) is immediately obtained by replacing $x$ in (2.15) with 1x1-x. We have
@@
By the change of variable $s=((1-x)t/x)^{\alpha}$ and [11, Equation 12, page 329], the right-hand side of (2.17) becomes
@@
as $n\to\infty$ for some $\delta>0$. (2.13) follows from (1.3), (2.19) and Lemma 2.7. Furthermore, Lemma 2.3 implies that
@@
as $n\to\infty$ for some $\delta>0$, by using (2.19) and (2.20). Moreover, (2.21), (2.22) and Lemma 2.7 imply that the modulus of the sum of the third and fourth terms on the right-hand side of (2.6) is dominated by
@@
which is $(|c|^{m}+1)O(e^{-\delta n})$ as well. Thus, (2.14) holds. Note that these items are regarded as 0 if $\alpha\in\mathbb{N}$. ∎
We often use Proposition 2.6 (2) with $c=-\alpha nx$ in the rest of this paper. When $c=-\alpha nx$, the second term on the right-hand side of (2.14) is $O(e^{-\delta n})$ with possibly different $\delta>0$. A direct calculation also gives
@@
Let $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is $\mu_{\alpha,x}^{(n)}$. Then, for some $\delta>0$, we have
@@
Applying Proposition 2.6 (2) with $c=0$ and $m=1$, we have
@@
as $n\to\infty$. This and Proposition 2.6 (1) imply (2.24). We also apply (2.23) and Proposition 2.6 (2) with $c=-\alpha nx$ and $m=2$ to obtain
@@
2.3 Uniform bound of $Z_{\alpha,x}^{(n)}$
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. We establish a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to $x$ and $n$, and again we note that $Z_{1,x}^{(n)}=1$.